OpenStudy (anonymous):

1/2 X + 6 = 52 92 46 26 104

OpenStudy (anonymous):

92

OpenStudy (anonymous):

there are 2 ways to do this... 1/2x + 6 = 52 -- subtract 6 from both sides 1/2x = 52 - 6 1/2x = 46 -- divide both sides by 1/2 to get x by itself (1/2x)/(1/2) = 46/(1/2) x = 46 * 2/1 x = 92 or this way, if you do not want to deal with fractions 1/2x + 6 = 52 -- multiply the entire equation by the LCD which is 2 x + 12 = 104 -- subtract 12 from both sides x = 104 - 12 x = 92